OpenStudy (anonymous):

Solve. Solve. x – 9 = –15 A. –6 B. 24 C. 6 D. –24 x – 9 = –15 A. –6 B. 24 C. 6 D. –24

OpenStudy (cheerleader):

a

OpenStudy (compassionate):

Cheerleader, it's best to lead the user to an answer. @Emo<3girl , you need to add +9 to both sides of the equation. This is how we solve equations. x - 9 = -16 x + 9 - 9 = -16 + 9 x = -6